The Missouri River is 149 miles longer than double the Snake River. The Potomac River is 796 miles shorter than the Snake River. The combined lengths of these 3 rivers is 3,685 miles. Find the length of each.
Answer by mananth(16946)

The Missouri River is 149 miles longer than double the Snake River. The Potomac River is 796 miles shorter than the Snake River. The combined lengths of these 3 rivers is 3,685 miles. Find the length of each.
snake river length be x miles
missouri river length will be x+149
Potomac river length = x-796
x+x+149+x-796=3685
3x= 3685=796-149
3x=4332
x=1444 miles length of snake river
1444+149= 1593 length of missouri river
Potomac river will be 1444-796=1648 miles